OpenStudy (anonymous):

how many solutions does the equation have? 6-a = 1/3(9a-18) a. one solution b. no solution c. infinite solution

TheSmartOne (thesmartone):

@anx Please refrain from giving direct answers... You need to guide them to it. And the asker should participate. http://openstudy.com/code-of-conduct

OpenStudy (anx):

I'll rephrase, sorry mate. 6 - a = distribute the (1/3) so 6 - a = (3a - 6) use algebra from this point to find a.

TheSmartOne (thesmartone):

Also note \(\Huge\color{red}{x}\color{blue}{(y+z)}=\color{green}{xy+xz}\)